OpenStudy (anonymous):

Solve the pair of simultaneous equations. y=8-x 2x square+ xy=-16

OpenStudy (anonymous):

2x^2 + x(8-x) = -16 2x^2 +8x - x^2 = -16 x^2 + 8x + 16 = 0 (x + 4)((x + 4) = 0 x = -4 Back to first equation y = 8 - - 4 y = 12

OpenStudy (anonymous):

what does the sign ^ mean

OpenStudy (anonymous):

it means that number is squared

OpenStudy (anonymous):

i dnt get it @gshrand im not used to this sign ^

OpenStudy (anonymous):

so instead of saying 2x squared im saying 2x^2 If i wanted to say 2x cubed i would say 2x^3

OpenStudy (anonymous):

oooook

OpenStudy (anonymous):

thank you for your help

OpenStudy (anonymous):

No problem!